Of the water lilies in the pond, 43% are yellow. The others are white. A frog randomly jumps onto a lily. Describe the complement of the frog landing on a yellow lily and find its probability. what is the answer

Answer by Edwin McCravy(20081)

The complement of a given event = the event that the given event does
not happen. 

The complement of the event of the frog landing on a yellow lily is
the event of the frog landing on a white lily instead.

P(complement of a given event) = 1 - P(given event)

P(frog landing on a white lily) = 1 - P(frog landing on a yellow lily)

P(complement event) = 1 - 0.43 = 0.57 or 57%

Edwin